OpenStudy (anonymous):

What is the equation of the line that passes through the points (−3, 2) and (−5, 8)?

Mehek (mehek14):

do you know how to find the slope?

OpenStudy (welshfella):

first find the slope of the line do you know how to do that?

OpenStudy (anonymous):

y2 - y1/ x2-x1

OpenStudy (madhu.mukherjee.946):

@kyle55 my formula

OpenStudy (anonymous):

thank you :)

OpenStudy (madhu.mukherjee.946):

now just do it

Mehek (mehek14):

correct so plug in your ordered pairs

OpenStudy (madhu.mukherjee.946):

:)

OpenStudy (anonymous):

8-2/ -5 + 3

Mehek (mehek14):

good so far

OpenStudy (anonymous):

6/-2 -3

Mehek (mehek14):

ok so m = -3

Mehek (mehek14):

\(\bf{y=\color{blue}{-3}x+\color{red}{b}}\)

Mehek (mehek14):

choose one of the ordered pairs you have

OpenStudy (anonymous):

y= -3(-5) + 8

OpenStudy (anonymous):

7= 15 + 8

OpenStudy (anonymous):

Y**

Mehek (mehek14):

no remember 8 goes instead of y

OpenStudy (anonymous):

oh so 8= 15 + b

OpenStudy (anonymous):

-7= b

Mehek (mehek14):

\(\bf{8=\color{blue}{-3}(-5)+\color{red}{b}}\)

Mehek (mehek14):

yes correct b = -7

Mehek (mehek14):

so you have the slope -3 and y-intercept -7

OpenStudy (anonymous):

what would the formula be?

OpenStudy (anonymous):

equation*

Mehek (mehek14):

\(\bf{y=\color{blue}{m}x+\color{red}{b}}\) slope = m = -3 y-intercept = b = -7

OpenStudy (anonymous):

3x + y = − 7

OpenStudy (anonymous):

?

OpenStudy (anonymous):

3x − y = −11 3x + y = − 7 . 3x + y = 5 x + 3y = 3

Mehek (mehek14):

\(\bf{y=\color{blue}{-3}x-\color{red}{7}}\)

OpenStudy (anonymous):

answers

Mehek (mehek14):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @kyle55 3x + y = − 7 \(\color{#0cbb34}{\text{End of Quote}}\) correct ^_^

OpenStudy (anonymous):

thank you :)

Mehek (mehek14):

yw ^_^